# Utilizing Large Language Models to Optimize the Detection and Explainability of Phishing Websites

### v1

@@ 3.2. Training data
To establish a ground-truth dataset for training our classifier, we utilized Lin at el’s PhishPedia dataset (Lin et al., 2021), which includes 30K phishing and 30K benign websites. PhishPedia is not only the largest publicly available phishing dataset but is also notable for its reliability, having been used in several prior literature (Liu et al., 2022; Guo et al., 2021). Each entry in this dataset typically includes associated metadata, such as HTML source code, screenshots, and OCR text. The phishing URLs were originally obtained from OpenPhish (Openphish, [n. d.]), a widely recognized antiphishing blocklist. These URLs underwent preprocessing to eliminate inactive sites and false positives, and the researchers also implemented strategies to counteract evasive tactics like cloaking and conducted manual verifications to correct any inaccuracies in target brand representations. Conversely, the benign URLs were sourced from the Alexa Rankings list (ale, 2020), a (now defunct) online database that ranked websites based on popularity. For the purposes of training our model, we only required the HTML source file, such that it can be parsed into a collection of relevant segments as previously detailed in Section 3.1. Out of the 30K phishing entries, 22,419 had an HTML file, whereas the same number for benign URLs was 26K. We further analyzed this dataset beyond targeted organizations (Lin et al., 2021), to obtain the distribution of HTML tags, libraries and frameworks used, and JavaScript APIs and function calls. Table 1 illustrates the distribution of HTML tags in both our phishing and benign websites. Notably, phishing websites contain fewer tags than benign websites. This is because phishing websites primarily focus on capturing user credentials, whereas benign websites need to offer a wide range of functionalities, requiring a more extensive and complex HTML structure. To identify the technology stack utilized by the phishing and benign websites, we used Wappalyzer (Chorsley, 2024), a tool that uses various heuristics and pattern-matching techniques to identify the libraries and frameworks implemented in websites. Wappalyzer has been frequently used in prior literature to characterize the technological stack of websites (He et al., 2017) Figures 2A. and Figure 2B. show the distribution of these frameworks across phishing and benign websites in the PhishPedia dataset. It is evident from the figures that phishing websites exhibit less diversity in the frameworks used compared to benign websites. This is because phishing websites only need the frameworks necessary to carry out their attacks, while benign websites use a variety of frameworks to support the myriad of features required by their users. We also focus on the distribution of JavaScript function calls used by phishing websites by utilizing Esprima (esp, [n. d.]), an ECMAScript (JavaScript) parser that generates an Abstract Syntax Tree (AST), allowing us to identify JS function calls and categorize the usage of specific APIs. We include this illustration in Fig 2C. Prior work (Fass et al., 2018) noted how Esprima has been reliable for this purpose.
Finally, we are interested in identifying whether PhishLang can detect evasive phishing attacks, although our model cannot categorize the types of evasion directly. Therefore, we applied separate heuristics, derived from specific studies (Oest et al., 2020a; Team, 2022; Auth0, 2021; Liang et al., 2016; Ventures, 2019), to classify the phishing URLs in the Phishpedia dataset into five categories: i) Regular phishing attacks (18,401), ii) Behavioral JS Evasions (715), iii) Clickjacking attacks (1,726), iv) Exploiting DOM (536), and v) Text encoding attacks (1,041). Behavioral JS Evasions, as highlighted by Oest et al.(Oest et al., 2020a), include JS-based redirection triggered by user interactions, mouse movement detections, and popup interactions, exploiting weaknesses in several prominent blocklists. In contrast, DOM Manipulations involve altering the HTML Document Object Model (DOM) to evade detection, such as dynamically generating phishing content, hiding malicious elements with CSS, and modifying the DOM structure post-page load. Liang et al. (Liang et al., 2016) identified features of these tactics that bypass Google’s phishing page filter (GPPF), which are also transferable to other detection models. Text encoding attacks include obfuscating the character encoding, nested encoding, encoding to obfuscate JavaScript (JS), and white space obfuscation, hiding malicious code within seemingly harmless text by manipulating how the text is interpreted by different browsers.
3.3. Training and inference
During the training phase, we propose a sliding window approach, where each website sample, represented as parsed HTML content, is processed using this approach. The window size, $W$, is set to the maximum token count that the LLM can handle. However, to ensure coherence and completeness, the window is adjusted to include only complete tags. This adjustment is crucial as it prevents the misinterpretation of partial data. Each window of content inherits the label of the entire website. Despite the possibility that some sections that are benign are incorrectly assigned the phishing label due to belonging to a phishing website a window, given that the majority of the phishing websites were shorter, with most having fewer than 900 tokens (compared to benign sites having a media of 2,415 tokens), it is unlikely for a chunk not to have significant phishing features. The model is then trained on these chunks, $w_{i}$, defined as:
$w_{i}=\text{content}[i\times S+1:\min((i\times S+W),T)]$
Where $S$ is the step size or stride of the sliding window. It determines how much the window moves forward after each iteration, and $T$ is the total token count of the website’s content. It represents the length of the content in terms of tokens, which is used to determine the boundaries for the sliding window approach. Here $i$ ranges from 0 to $\lceil(T-W)/S\rceil$, ensuring full coverage of the content, and each window $w_{i}$ is assigned the label $y$, corresponding to the entire website’s classification (either phishing or benign). We use 70% of our ground truth data for training the model and reserve 30% for testing. The model is trained using 5-fold cross-validation, with each fold undergoing 10 epochs of training.
The inference process mirrors the training setup. The model analyzes the parsed content of a website by breaking it down into chunks using the same sliding window technique. Each chunk is evaluated independently for signs of phishing. If any chunk is predicted as phishing, the entire website is classified accordingly:
Website $\text{Website Classification}=\begin{cases}\text{Phishing}&\text{if }\exists\,p_{i}=\text{Phishing}\\ \text{Benign}&\text{otherwise}\end{cases}$
To counteract evasion techniques where attackers might pad websites with non-functional elements to dilute phishing content across chunks (Panum et al., 2020), a dynamic merging strategy is employed. If all chunks are initially predicted as benign, they are progressively merged and re-evaluated. This merging starts with the first two chunks, and if the combined chunk remains benign, it is further merged with the next chunk, continuing until a definitive phishing signal is detected or all chunks have been evaluated:
$p_{merge}=\text{Model}(w_{1}+w_{2}+\ldots+w_{k})$
Final $\text{Final Classification}=\begin{cases}\text{Phishing}&\text{if }p_{merge}=\text{Phishing}\\ \text{Benign}&\text{otherwise}\end{cases}$
The sliding window method to break down the parsed website content into multiple chunks for both training and inference is essential for several reasons. Firstly, it circumvents the token limitations of the LLM models, as most of them can process only a finite number of tokens at a time. This segmentation maintains contextual integrity, ensuring that each portion of a webpage is analyzed within a coherent context. Moreover, this method enhances detection sensitivity and reduces false negatives in scenarios where only a small part of a webpage is malicious; thus, training on the entire page might dilute the malicious signals. Logistically, processing smaller chunks of data is also more efficient on multi-core processors and GPUs, which these models are trained on, significantly speeding up training and inference times.
3.4. Choosing the optimum model
@@
To choose the language model most suitable for building our classifier, we evaluated and compared eight language models that are considered to provide good performance in binary classification tasks (the number of parameters for the models in parenthesis): DistilBERT(66M) (Sanh et al., 2019), TinyBERT(15M) (Jiao et al., 2019), DeBERTA-base (He et al., 2020), FastText (10M) (Grave et al., 2018), and GPT-2 (117M) (Radford et al., 2019), Llama (7B), Bloom and T5. Our aim was to choose the model which provides the best trade-off between performance, speed, and memory usage. We trained each of these models on the training split (70%) of our ground truth dataset and tested them on our test split (30%).
@@
The performance of these models is illustrated in Table 2. We found Llama 2 to have the best performance out of all the models tested, with an F1 score of 0.96. However, its median inference time of  34s is impractical for real-world usage where a phishing detection tool would need to process hundreds of thousands of URLs every day, and it has the highest median memory usage at 4,873.47MB. While the other two LLMs, Bloom and T5, had lower median inference times of 7.49 and 12.40 seconds respectively, they also performed worse (F1 scores of 0.88 and 0.74, respectively) and had high memory requirements (7351.89MB for Bloom and 1,279.49MB for T5). GPT-2 had the lowest median inference time (1.81 seconds) but also had the worst performance among the LLMs (F1 0.68) and required 922.41MB of memory.
Moreover, all four LLMs required inference over GPUs (while they can be used for inference using the CPU, the prediction times will be much slower), and consequently required a large amount of VRAM for said inference per sample. Requiring GPU to provide good inference time is not ideal in a practical setting, as most consumer-end devices where the inference needs to occur (such as portable/mobile systems) might not have GPUs.
@@
When considering the Small Language Models (SLMs), DistilBERT had the highest accuracy level at 0.94, with both precision and recall impressively scoring 0.94. While not the quickest, its prediction times were reasonably efficient, with a median of 0.85 seconds per prediction and a memory usage of 502.17MB. In comparison, DeBERTa-base delivered lower precision and recall also fell short in time efficiency, and had a higher memory usage of 1,341.39MB. FastText, though more compact with a memory usage of 201.88MB, suffered from reduced accuracy and median prediction times of 1.43 seconds. Meanwhile, TinyBERT, the smallest model tested, offered quicker results at 0.78 seconds but at the cost of lower performance metrics and required 495.15MB of memory. All SLMs could provide inference over CPUs.
Also, due to the unavailability of commercial LLMs such as ChatGPT and Claude as local implementations, fine-tuning and testing on these models would incur significant costs for training and, especially for inference, given the millions of URLs an anti-phishing tool needs to process daily. However, we compare ChatGPT (3.5 Turbo) and GPT 4 with our model, and other ML-based phishing detection tools in Section 4.
Thus, considering all eight language models, based on a combination of high precision and recall, fast prediction times, and efficient memory usage, DistilBERT was the most suitable choice for our framework and was thus chosen as the final language model for our PhishLang framework. Recent studies have also found that foundation models (and their derivatives like DistilBERT) perform well for binary classification tasks (Kaliyar et al., 2021; Hu et al., 2023).
@@ 4.1. Identifying new URLs
We detected 17,396 regular, 3,159 JavaScript evasion, 3,349 Clickjacking, 1,057 DOM, and 835 Text encoding attacks. Some evasions might not have been identified and could have been classified as regular attacks. To verify the accuracy of the model, two coders, experienced in computer security concepts, manually assessed 2.5k websites flagged as phishing, ensuring a representative sample of each attack type. They first checked if the websites impersonated any of the 409 brands identified as common phishing targets by OpenPhish in August 2022 (OpenPhish, 2022), followed by examining if the sites solicited sensitive information or triggered suspicious downloads. Downloads with four or more detections on VirusTotal were deemed malicious, based on previous findings (Peng et al., 2019). The inter-rater reliability (Cohen’s kappa) was 0.82, indicating high agreement and any disagreements were mutually resolved. Ultimately, 2,388 sites were confirmed as phishing, demonstrating the model’s accuracy of approximately 94% for regular attacks and an average detection of 91% for the four evasive categories. We break down the model’s performance against each type of attack in Table 3. It is evident that PhishLang is capable of detecting a range of evasive phishing attacks.
@@
4.2. Comparison with other ML models
We compared the performance of PhishLang with other popular open-source ML-based phishing detection models. Two of these models rely on the visual features of the website: VisualPhishNet (Abdelnabi et al., 2020) and PhishIntention (Liu et al., 2022), one relies on both the URL string and HTML representation of the website: StackModel (Li et al., 2019), and one that relies on the semantic representation of the URL string only: URLNet (Le et al., 2018). Each of these models was tested on the 5K samples, which we had labeled earlier. We also include the commercially available models of ChatGPT (GPT 3.5T and GPT 4). Since the focus of PhishLang is not only on prediction performance but also on the overhead of the models, which is crucial for practical implementation, we add two metrics - Median Artifact size, i.e., the amount of data (source code, URL, image, etc. based on the model) that needs to be provided to the model for prediction, and Median Prediction Time, which is the median time required for the respective model to provide the prediction label. Also, since GPT 3.5T and 4 are general LLM models that can predict any text content, we pass both the full HTML and the parsed versions as artifacts for evaluation. Table 4 illustrates the performance of each of the models. We note that PhishIntention, despite showing the best performance among all the models, requires a relatively large artifact size of 348KB and has a longer prediction time of 10.7 seconds. PhishLang, on the other hand, shows strong performance metrics with the advantage of a significantly smaller artifact size (7KB) and faster prediction time (0.9 seconds), indicating a balance between high performance and operational efficiency. Visual PhishNet and StackModel show moderate performance in all metrics, with the former requiring a much larger artifact size and longer prediction time compared to StackModel. Finally, URLNet has the lowest performance metrics but benefits from the smallest artifact size (0.2KB) and shortest prediction time (0.7 seconds). In the case of the ChatGPT models, we found that GPT 3.5 was only able to process 822 phishing + 1031 benign samples when the full HTML was given, while GPT 4 was able to process 1,469 phishing + 2,147 benign samples due to the token size limit. We find that both models perform very well against the threats and are almost on par with both PhishLang and PhishIntention. Note that these GPT models are general models and were not specifically fine-tuned using phishing training data. Thus, it is possible that it can outperform our model if fine-tuned on appropriate data. However, restricting GPT models for commercial use and the overall expense required for training them could make them unfeasible for practical implementation.
Table 4. Comparison with other ML-based detection methods
Model Accuracy Precision Recall F1 Score Size Time
PhishLang 0.94 0.93 0.95 0.94 7KB 0.9
PhishIntention 0.96 0.94 0.96 0.95 348KB 10.7
Visual PhishNet 0.85 0.83 0.86 0.85 217KB 4.9
URLNet 0.73 0.72 0.74 0.73 0.2KB 0.7
StackModel 0.83 0.85 0.81 0.82 1KB 1.9
GPT 3.5T (HTML)* 0.87 0.88 0.86 0.87 42KB 2.1
GPT 4 (HTML)* 0.92 0.93 0.91 0.92 42KB 2.8
GPT 3.5T (Parsed) 0.90 0.90 0.88 0.88 7KB 1.6
GPT 4 (Parsed) 0.94 0.93 0.93 0.93 7KB 1.4
@@ 4.3. Evaluating Commercial Phishing Countermeasures
In this section, we evaluate the effectiveness of different phishing countermeasures against websites identified by PhishLang. These include two commercial browser protection tools—Google Safe Browsing and Microsoft SmartScreen—and two open-source phishing blocklists, PhishTank (phi, 2020b) and OpenPhish (Openphish, [n. d.]). Safe Browsing is the default protection in Google Chrome, Mozilla Firefox, and Safari, while SmartScreen powers Microsoft Edge, collectively protecting nearly 95% of all internet users (Stats, [n. d.]). We assess how many phishing websites identified by PhishLang are covered by these tools and whether our reporting helps detect the initially missed websites. PhishTank and OpenPhish, community-driven blocklists, are used by several commercial antiphishing tools (phi, 2020a; Oest et al., 2020b).
PhishLang checks each identified phishing URL against these tools and blocklists in real time. If a URL is not covered, it is reported, and we monitor its detection status for up to a week. Considering ethical implications, we report all 26k identified websites immediately to ensure user protection, despite our 97% accuracy rate suggesting a potential 3% false positives being reported. This trade-off is acceptable given the typically higher noise in blocklist data (Saha Roy et al., 2023; Roy et al., 2021; Oest et al., 2020b).
@@
Browser protection tools employ heuristic and ML-based models within the browser environment to detect phishing pages (Liang et al., 2016). Using Selenium (documentation, [n. d.]), we automate opening each website in a browser and take screenshots to confirm detection by these tools. Table 5 illustrates the initial coverage of these tools and subsequent coverage after we had reported the URLs that they had failed to detect.
All tools initially have low coverage, with Google Safe Browsing outperforming SmartScreen. Reporting undetected URLs significantly improved detection rates, especially for evasive phishing attacks. For instance, Google Safe Browsing’s detection of JS evasion attacks increased from 35% to 73%, and SmartScreen’s detection of DOM-based attacks rose from 18% to 86%. Regular phishing URL detection also improved significantly. Despite these improvements, it is important to note that these tools generally maintain high detection rates and low false positives (Oest et al., 2020b). Our model thus helps close detection gaps, particularly for new and evasive attacks, and accelerates the integration of new threat intelligence.
PhishTank and OpenPhish have different responses to reported URLs. PhishTank lists reported URLs immediately, requiring community verification before labeling them as phishing or benign, and we measure the time taken between our reporting and subsequent labeling. OpenPhish verifies URLs before inclusion in the blocklist, and we measure the time until URLs appear on the blocklist. For hosting domains, we track how long it takes for reported websites to become inactive. Table 5 again illustrates the performance of these entities, showing significant detection rate improvements following our reports, especially for evasive attacks, indicating that PhishLang enhances these countermeasures’ effectiveness.

### v2

@@ 3.2.1. Groundtruth
To establish a ground-truth dataset for training our classifier, we utilized Lin at el’s PhishPedia dataset (Lin et al., 2021), which includes 30K phishing and 30K benign websites. PhishPedia is not only the largest publicly available phishing dataset but is also notable for its reliability, having been used in several prior literature (Liu et al., 2022; Guo et al., 2021). Each entry in this dataset typically includes associated metadata, such as HTML source code, screenshots, and OCR text. The phishing URLs were originally obtained from OpenPhish (Openphish, [n. d.]), a widely recognized antiphishing blocklist. These URLs underwent preprocessing to eliminate inactive sites and false positives, and the researchers also implemented strategies to counteract evasive tactics like cloaking and conducted manual verifications to correct any inaccuracies in target brand representations. Conversely, the benign URLs were sourced from the Alexa Rankings list (ale, 2020), a (now defunct) online database that ranked websites based on popularity. For the purposes of training our model, we only required the HTML source file, such that it can be parsed into a collection of relevant segments as previously detailed in Section 3.1. Out of the 30K phishing entries, 22,419 had an HTML file, whereas the same number for benign URLs was 26K. We further analyzed this dataset beyond targeted organizations (Lin et al., 2021), to obtain the distribution of HTML tags, libraries and frameworks used, and JavaScript APIs and function calls. Table 1 illustrates the distribution of HTML tags in both our phishing and benign websites. Notably, phishing websites contain fewer tags than benign websites. This is because phishing websites primarily focus on capturing user credentials, whereas benign websites need to offer a wide range of functionalities, requiring a more extensive and complex HTML structure. We also used Wappalyzer (Chorsley, 2024) and Esprima (esp, [n. d.]) to identify the frameworks and JS functions used by and present respectively in these websites. For brevity, we include this analysis in Section 9.2 of the Appendix. Finally, we are interested in identifying whether PhishLang can detect evasive phishing attacks, although our model cannot categorize the types of evasion directly. Therefore, we applied separate heuristics, derived from specific studies (Oest et al., 2020a; Team, 2022; Auth0, 2021; Liang et al., 2016; Ventures, 2019), to classify the phishing URLs in the Phishpedia dataset into five categories: i) Regular phishing attacks (18,401), ii) Behavioral JS Evasions (715), iii) Clickjacking attacks (1,726), iv) Exploiting DOM (536), and v) Text encoding attacks (1,041). Behavioral JS Evasions, as highlighted by Oest et al.(Oest et al., 2020a), include JS-based redirection triggered by user interactions, mouse movement detections, and popup interactions, exploiting weaknesses in several prominent blocklists. In contrast, DOM Manipulations involve altering the HTML Document Object Model (DOM) to evade detection, such as dynamically generating phishing content, hiding malicious elements with CSS, and modifying the DOM structure post-page load. Liang et al. (Liang et al., 2016) identified features of these tactics that bypass Google’s phishing page filter (GPPF), which are also transferable to other detection models. Text encoding attacks include obfuscating the character encoding, nested encoding, encoding to obfuscate JavaScript (JS), and white space obfuscation, hiding malicious code within seemingly harmless text by manipulating how the text is interpreted by different browsers. We use 70% of our ground truth data for training the model and reserve 30% for testing. The model is trained using 5-fold cross-validation, with each fold undergoing 10 epochs of training.
3.2.2. Choosing the optimum model
@@
To choose the language model most suitable for building our classifier, we evaluated and compared nine language models that are considered to provide good performance in binary classification tasks (the number of parameters for the models in parenthesis): DistilBERT(66M) (Sanh et al., 2019), TinyBERT(15M) (Jiao et al., 2019), DeBERTA-base (He et al., 2020), MobileBERT (25M) (Sun et al., 2020), FastText (10M) (Grave et al., 2018), and GPT-2 (117M) (Radford et al., 2019), Llama (7B), Bloom and T5. Our aim was to choose the model which provides the best trade-off between performance, speed, and memory usage. We trained each of these models on the training split (70%) of our ground truth dataset and tested them on our test split (30%).
@@
The performance of these models is illustrated in Table 2. We found Llama 2 to have the best performance out of all the models tested, with an F1 score of 0.96. However, its median inference time of  34s is impractical for real-world usage where a phishing detection tool would need to process hundreds of thousands of URLs every day, and it has the highest median memory usage at 4,873.47MB. While the other two LLMs, Bloom and T5, had lower median inference times of 7.49 and 12.40 seconds respectively, they also performed worse (F1 scores of 0.88 and 0.74, respectively) and had high memory requirements (7351.89MB for Bloom and 1,279.49MB for T5). GPT-2 had the lowest median inference time (1.81 seconds) but also had the worst performance among the LLMs (F1 0.68) and required 922.41MB of memory. Moreover, all four LLMs required inference over GPUs (while they can be used for inference using the CPU, the prediction times will be much slower), and consequently required a large amount of VRAM for said inference per sample. Requiring GPU to provide good inference time is not ideal in a practical setting, as most consumer-end devices where the inference needs to occur (such as portable/mobile systems) might not have GPUs.
@@
When considering the Small Language Models (SLMs), MobileBERT had the highest accuracy level at 0.96, with both precision and recall scoring 0.95 and 0.96, respectively. It also demonstrated the fastest prediction times among the SLMs, with a median of 0.39 seconds per prediction, and required the least memory usage at 74MB among all models tested. These attributes make MobileBERT highly suitable for real-time phishing detection, balancing high performance with low computational resource requirements, which is ideal for deployment on devices with limited hardware capabilities, such as mobile and portable systems. Its efficient use of DRAM instead of requiring a GPU makes it even more attractive for widespread, scalable deployment. In comparison, DistilBERT achieved an accuracy of 0.94, with both precision and recall scoring 0.94. While not the quickest, its prediction times were reasonably efficient, with a median of 0.85 seconds per prediction and significantly higher memory usage of 502MB compared to MobileBERT. DeBERTa-base, another SLM, delivered lower precision and recall, also fell short in time efficiency, and had a higher memory usage of 1,341MB. FastText, though more compact with a memory usage of 201MB, suffered from reduced accuracy (0.62) and had a median prediction time of 1.43 seconds. Meanwhile, TinyBERT, the smallest model tested, offered quicker results at 0.78 seconds but at the cost of lower performance metrics, requiring 495MB of memory. All SLMs could provide inference over CPUs, making them suitable for deployment on various platforms. Also, due to the unavailability of commercial LLMs such as ChatGPT and Claude as local implementations, fine-tuning and testing on these models would incur significant costs for training and, especially for inference, given the millions of URLs an anti-phishing tool needs to process daily. However, we compare ChatGPT (3.5 Turbo) and GPT 4 with our model, and other ML-based phishing detection tools in Section 8 in the Appendix. Thus, considering all the tested language models, based on a combination of high precision and recall, fast prediction times, and efficient memory usage, MobileBERT was the most suitable choice for our framework and was thus chosen as the final language model for our PhishLang framework.
@@ 4.1. Identifying new URLs
We detected 17,396 regular, 3,159 JavaScript evasion, 3,349 Clickjacking, 1,057 DOM, and 835 Text encoding attacks. Some evasions might not have been identified and could have been classified as regular attacks. To verify the accuracy of the model, two coders, experienced in computer security concepts, manually assessed 2.5k websites flagged as phishing, ensuring a representative sample of each attack type. They first checked if the websites impersonated any of the 409 brands identified as common phishing targets by OpenPhish in August 2022 (OpenPhish, 2022), followed by examining if the sites solicited sensitive information or triggered suspicious downloads. Downloads with four or more detections on VirusTotal were deemed malicious, based on previous findings (Peng et al., 2019). The inter-rater reliability (Cohen’s kappa) was 0.82, indicating high agreement and any disagreements were mutually resolved. Ultimately, 2,388 sites were confirmed as phishing, demonstrating the model’s accuracy of approximately 94% for regular attacks and an average detection of 91% for the four evasive categories. We break down the model’s performance against each type of attack in Table 3. It is evident that PhishLang is capable of detecting a range of evasive attacks.
@@ 4.2. Evaluating Commercial Phishing Countermeasures
In this section, we evaluate the effectiveness of different phishing countermeasures against websites identified by PhishLang. These include two commercial browser protection tools—Google Safe Browsing and Microsoft SmartScreen—and two open-source phishing blocklists, PhishTank (phi, 2020b) and OpenPhish (Openphish, [n. d.]). Safe Browsing is the default protection in Google Chrome, Mozilla Firefox, and Safari, while SmartScreen powers Microsoft Edge, collectively protecting nearly 95% of all internet users (Stats, [n. d.]). We assess how many phishing websites identified by PhishLang are covered by these tools and whether our reporting helps detect the initially missed websites. PhishTank and OpenPhish, community-driven blocklists, are used by several commercial antiphishing tools (phi, 2020a; Oest et al., 2020b). PhishLang checks each identified phishing URL against these tools and blocklists in real time. If a URL is not covered, it is reported, and we monitor its detection status for up to a week. Considering ethical implications, we report all 26k identified websites immediately to ensure user protection, despite our 97% accuracy rate suggesting a potential 3% false positives being reported. This trade-off is acceptable given the typically higher noise in blocklist data (Saha Roy et al., 2023; Roy et al., 2021; Oest et al., 2020b).
@@
. Browser protection tools employ heuristic and ML-based models within the browser environment to detect phishing pages (Liang et al., 2016). Using Selenium (documentation, [n. d.]), we automate opening each website in a browser and take screenshots to confirm detection by these tools. Table 4 illustrates the initial coverage of these tools and subsequent coverage after we had reported the URLs that they had failed to detect. All tools initially have low coverage, with Google Safe Browsing outperforming SmartScreen. Reporting undetected URLs significantly improved detection rates, especially for evasive phishing attacks. For instance, Google Safe Browsing’s detection of JS evasion attacks increased from 35% to 73%, and SmartScreen’s detection of DOM-based attacks rose from 18% to 86%. Regular phishing URL detection also improved significantly. Despite these improvements, it is important to note that these tools generally maintain high detection rates and low false positives (Oest et al., 2020b). Our model thus helps close detection gaps, particularly for new and evasive attacks, and accelerates the integration of new threat intelligence.
PhishTank and OpenPhish have different responses to reported URLs. PhishTank lists reported URLs immediately, requiring community verification before labeling them as phishing or benign, and we measure the time taken between our reporting and subsequent labeling. OpenPhish verifies URLs before inclusion in the blocklist, and we measure the time until URLs appear on the blocklist. For hosting domains, we track how long it takes for reported websites to become inactive. Table 4 shows the performance of these entities, showing significant detection rate improvements following our reports, especially for evasive attacks, indicating that PhishLang enhances these countermeasures’ effectiveness.